A machine costing $209,800 with a four-year life and an estimated $15,000 salvage value is installed in Luther Company’s factory on January 1. The factory manager estimates the machine will produce 487,000 units of product during its life. It actually produces the following units: 122,600 in Year 1, 123,300 in Year 2, 120,800 in Year 3, 130,300 in Year 4. The total number of units produced by the end of Year 4 exceeds the original estimate—this difference was not predicted. (The machine cannot be depreciated below its estimated salvage value.)CULORUMLUSUNULIC W. hol predicted (the machine cannot be depreciated below its estimated salvage value.) Required: 1.42 pointestimated salvage value) Required: 142 polres Compute depreciation for each year and total depreciation of all years combined
Answer #1

Calculation of Depreciation:

1. Straight Line Method:

Depreciation = (Cost of the asset - Residual value) / Useful life of the asset

Cost of the Asset = $209,800

Salvage value = $15,000

Useful life of the asset = 4 Years

Depreciation = ($209,800 - $15,000) / 4 Years = $194,800 / 4 Years = $48,700

Year Opening Value Depreciation Closing value
1 $209,800 $48,700 $161,400
2 $161,400 $48,700 $112,400
3 $112,400 $48,700 $63,700
4 $63,700 $48,700 $15,000

2. Units of Production Method:\

Depreciation =

Annual Depreciation = Depreciable Value × (Units produced during the year / Estimated total production )

Depreciable Value = Cost of the asset - Salvage Value = $209,800 - $15,000 = $194,800

Total Units Produced during the life = 487,000 Units

Year Units Produced Depreciable Value Calculation Depreciation Expense
1 122,600 ($194,800 / 497,000) = 0.39 $194,800 * ( 122,600 / 497,000 ) $48,053
2 123,300 ($194,800 / 497,000) = 0.39 $194,800 * (123,300 / 497,000) $48,328
3 120,800 ($194,800 / 497,000) = 0.39 $194,800 * (120,800 / 497,000) $47,348
4 130,300 ($194,800 / 497,000) = 0.39 $194,800 * (130,300 / 497,000) $51,071
Total 497,000

3.Double Decling Method:

Depreciation = 2 X Cost of the asset X Depreciation rate

Cost of the Asset = $209,800

Depreciation rate = 100% / Useful Life = 100%/4 Years = 25%

Year Calculation Opening value Depreciation Closing value
1 $209,800*2*25% $209,800 $104,900 $104,900
2 $104,900*2*25% $104,900 $52,450 $52,450
3 $52,450*2*25% $52,450 $26,225 $26,225
4 $26,225*2*25% $26,225 $13,112 $13,113